Solve the problem.

Find the chi-square value  corresponding to a sample size of 23 and a confidence level of 90%.

16.013

33.924

1.69

12.338
Answer #1

For 90% confidence interval, α = 1 - 0.90 = 0.10,   α/2 = 0.05    
sample size = n = 23    
Degrees of freedom = df = n - 1 = 23 - 1 = 22    
df = 22    
We find Right value Chi Square using Excel Function CHISQ.INV.RT    
χ² critical = CHISQ.INV.RT(0.05, 22)    
χ² critical = 33.9244